A ray of light in the air strikes a glass plate (n=1.5) at an incidence angle of 50°. determine the angle of transmission.

1
Expert's answer
2022-01-20T10:07:45-0500

"n=\frac{\sin\alpha}{\sin \beta},"

"\beta=\arcsin\frac{\sin\alpha}n=30.7°."